The process used by the Gourmet Food Company (Process 1) to produce dressings has annual fixed costs of $240,000 and variable costs of $0.50 per bottle. The company just entered into an agreement with a major national grocery store chain to sell its dressings. Sales volumes are expected to increase. Two new processes are being explored. Process 2 has a fixed cost of $320,000 per year and variable costs of $.30 per bottle. Process 3 has fixed costs of $400,000 per year and variable costs of $.25 per bottle.

If sales are expected to be 300,000 bottles, which process should be used?

a. Process 1

b. Process 2

c. Process 3
Answer #1

In this problem, we'll select the process with the lowest total cost. The total cost can be computed using the formula:

Total cost = Fixed cost + (Sales volume x Variable cost per unit)

We know that Sales volume = 300,000 bottles

For process 1,

Fixed cost = $240,000

Variable cost per unit = $0.50

Total cost = 240,000 + (300,000 x 0.50) = $390,000

For process 2,

Fixed cost = $320,000

Variable cost per unit = $0.30

Total cost = 320,000 + (300,000 x 0.30) = $410,000

For process 3,

Fixed cost = $400,000

Variable cost per unit = $0.25

Total cost = 400,000 + (300,000 x 0.25) = $475,000

Since the total cost of Process 1 is the lowest, We'll choose Process 1.